OpenStudy (anonymous):

Solve by factoring the GCF: 16^2-4x=0

OpenStudy (anonymous):

\[4x(4x-1)=0\] \[4x=0\iff x=0\] \[4x-1=0\iff 4x=1\iff x=\frac{1}{4}\]

myininaya (myininaya):

are you sure this is right?

OpenStudy (anonymous):

lol i read it as \[16x^2-4x=0\] silly me. back to work!

myininaya (myininaya):

its possible she meant 16x^2-4x=0